OpenStudy (anonymous):

Which of the following is the least profitable investment for a candy shop that earns $1 profit per pound of candy? Worker at $10 per hour, producing eight pounds of candy per hour Worker at $12 per hour, producing 16 pounds of candy per hour Machine with $5 per hour operating cost, producing 10 pounds of candy per hour Machine with $8 per hour operating cost, producing 14 pounds of candy per hour

OpenStudy (anonymous):

I think its Worker at $10 per hour, producing eight pounds of candy per hour

OpenStudy (anonymous):

Also I need help with this one: Which of the following is the least profitable investment for a game shop earning $2 profit from every game sold? Vehicle with $120 daily operating cost, delivering 80 games per day Inventory software with $10 daily operating cost, saving 30 games per day from theft Game cleaning system with $5 daily operating cost, increasing sales by 15 games per day Electronic billboard with $20 daily operating cost, increasing sales by 20 games per day

OpenStudy (anonymous):

For the second one, I think its Game cleaning system with $5 daily operating cost, increasing sales by 15 games per day

OpenStudy (radar):

I would look a little closer at the second one. Remember you are looking for the "least" profitable.

OpenStudy (anonymous):

So Machine with $5 per hour operating cost, producing 10 pounds of candy per hour?

OpenStudy (radar):

No, the second problem...

OpenStudy (radar):

You are correct for the first problem.

OpenStudy (radar):

What is profit from using the electronic billboard?

OpenStudy (anonymous):

$2

OpenStudy (anonymous):

Oh never mind $20

OpenStudy (radar):

compare that with the other options, the cleaning system which cost $5 daily brought in 15 extra sales at a profit of $2 per sale giving you a profit of $25.00 That is larger than the billboard, and the problem is looking for the least profitable.

OpenStudy (anonymous):

So it would probably be Inventory software with $10 daily operating cost, saving 30 games per day from theft

OpenStudy (radar):

The profit from the inventory software would be $50.00 which is even more than the billboard. Remember it is asking for the "least" profitable!

OpenStudy (anonymous):

so the only one left is the first option...but I thought it came out to $40

OpenStudy (radar):

It did, I asked you to look at the last option. What is the profit of the electronic billboard? You said $20, you are correct........Isn't it the least?

OpenStudy (anonymous):

Oh right...I'm not very good at this. Thanks for explaining though ^-^

OpenStudy (radar):

You're welcome, and good luck with your studies.

OpenStudy (anonymous):

With the video game one~ A. Net daily = (80)($2) - $120 = 160 - 120 = +$40 B. Net daily = (30)($2) - $10 = 60 = 10 = +$50 C. Net daily = (15)($2) - $5 = 30 - 5 = +$25 D. Net daily = (20)($2) - $20 = 40 - 20 = +$20